OpenStudy (anonymous):

What is the solution for this inequality? -4x < 28

OpenStudy (anonymous):

Divide -4 from both sides

OpenStudy (lynfran):

divide by -4 to both sides of the inequality and change the sign to >

OpenStudy (anonymous):

\( \huge \frac{-4x}{-4} = \frac{28}{-4} \) \( \huge x = \frac{28}{-4} \) What does x equal?

OpenStudy (lynfran):